OpenStudy (anonymous):

please help?! 350 is 70% of what number?

OpenStudy (hunus):

\[0.7*x=350\]

OpenStudy (tkhunny):

350 is all of 350 (100%) 350 is half of 700 (50%) 350 is one fourth of 1400 (25%)

OpenStudy (anonymous):

350 is 70% of what number (70/100)* X = 350 0.7 * X = 350 X = 350/0.7 X = 500.....this is the required answer